OpenStudy (anonymous):

r = -6 sin(theta) Part I: Multiply both sides of this equation by r and use the fact that to rewrite this equation in terms of x and y. Part II: Complete the square to produce the final equation. Part III: Using only your answer to part II, what shape does this graph make?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

did you do Part I ? what did you get?

OpenStudy (anonymous):

No I have no idea how to do. D:

OpenStudy (anonymous):

use the fact that : r^2 = x^2 + y^2 x = rcos(theta) y = rsin(theta) in part i, you should get x^2 + y^2 = -6rsin(theta) notice you can replace the rsin(theta) with y...

OpenStudy (thomas5267):

\[ \begin{align*} r&=-6\sin(\theta) \\ r^2&=r(-6\sin(\theta))&\text{Part I} \\ \frac{r^2}{-6}&=r\sin(\theta) \\ \frac{r^2}{-6}&=y \\ r^2&=-6y \\ x^2+y^2&=-6y \end{align*} \] Try complete the square and finish part ii and iii.

OpenStudy (anonymous):